# From #Dr00gtiktok to #harmreduction: Exploring Substance Use Hashtags on TikTok

### v1

@@ Methods
Given these patterns in how content creators use hashtags to frame and contextualize their videos, we next examined a key aspect of substance-related content: the inclusion of substances or their related paraphernalia themselves in the videos with respect to context and engagement. This analysis revealed distinct patterns in content distribution across three categories: no substances shown (n=260), substances shown but not used (n=68), and active substance use (n=23). Videos showing active substance use were predominantly focused on documentation for social media (approximately 40% of videos, Figure 5), with a stronger tendency towards lifestyle or vlog-style content. In contrast, recovery advocacy content was most prevalent in videos without substance visibility, while promotional content appeared most frequently in videos that showed but did not depict active use of substances (e.g., videos promoting nicotine products). Within videos showing substances, content predominantly featured alcohol, nicotine, or cannabis products, with only one video collage containing “hard drugs” such as Percocet, Adderall, and Xanax. A Kruskal-Wallis H (Kruskal and Wallis 1952) test was performed to compare the means in engagement metrics (views, likes, comments, and shares) between the substance presence categories. Notably, there were no statistically significant differences between groups. This suggests that user engagement on TikTok is driven more by content type and presentation than by the mere presence or use of substances.
Discussion
Social media’s impact on substance use behaviors has been a growing concern, with platforms often criticized for normalizing or promoting dangerous behaviors (Curtis et al. 2018; Moreno and Whitehill 2014; Lobstein et al. 2017; Rutherford et al. 2023; Romer and Moreno 2017). Our analysis of TikTok, however, revealed a more nuanced reality where recovery and harm reduction messaging dominate substance-related discourse. We found evidence of TikTok emerging as a significant platform for recovery support while maintaining its core identity as an entertainment platform.
The Power of Digital Recovery Narratives The most striking finding is the prevalence of recovery-focused content, which typically feature first-person narratives celebrating sobriety milestones or sharing recovery journeys. As Slade et al. (2021) note, such “digital recovery narratives” can increase knowledge of evidence-based treatment options, normalize addiction experiences, and encourage help-seeking behaviors. The impact of these narratives appears to be bidirectional - benefiting creators and viewers. For creators, sharing recovery stories can reinforce recovery-focused, collective identities and provide accountability through community support (Beck et al. 2023). For viewers, these narratives offer realistic models of recovery success, reduce feelings of isolation, and demonstrate that recovery is possible across different circumstances (Jacobson 2001; Kaiser et al. 2020; Corvini, Golomski, and Burns 2024).
Balancing Entertainment and Support in Health Communication The significant presence of humorous content (28.2%) alongside serious recovery narratives reflects TikTok’s unique ability to facilitate health communication through multiple approaches. Rather than undermining serious messages, humor appears to serve as an additional entry point for engaging with recovery-related content. The use of humor in substance-related content may serve important psychological functions, as research has shown that comedy can help individuals process difficult experiences (Martin 2002, 2003) and reduce stigma around mental health challenges (Corrigan et al. 2014). In the context of substance use and recovery on TikTok, videos that integrate humor appear to make the challenges around substance use more approachable while facilitating community connection through shared experiences. This balance between entertainment and support may be particularly effective for reaching younger audiences. However, this dual nature also presents challenges. While most humorous content avoids glorifying substance use some videos may inadvertently normalize concerning behaviors through comedy. This tension between engagement and responsible messaging represents a key challenge for platforms hosting health-related content.
Network Structure Facilitates Community Building The network analysis revealed a highly interconnected structure with small-world properties, enabling both focused discussions within communities and efficient information flow between them. Hashtags serve multiple functions: some act as content summarizers (e.g., “#harmreduction”, “#recovery”), while others serve as contextualizers (e.g., identity-based tags, occupation tags). Within this network structure, we observed the presence of algospeak in the form of coded or evasive hashtags (e.g., “#drvg”, “#drvgzx”, “#dr00gtiktok”) that are designed to evade content moderation. Our content analysis revealed that these coded hashtags were predominantly associated with satirical and relatable content, with a small subset appearing in recovery advocacy videos. This pattern suggests these alternative spellings may serve less to promote substance use and more to facilitate humor-based discussions and shared experiences around substance use, though this practice could complicate platforms’ efforts to effectively moderate substance-related content (Klug, Steen, and Yurechko 2023; Levine 2022). This dual functionality helps explain why content tagged with harm reduction and recovery-focused hashtags received substantially higher engagement despite appearing less frequently in our sample.
Looking Forward: Implications for Social Media Health Communication
Our findings have important implications for how social media platforms approach substance-related content, not just on TikTok but across all platforms featuring short-form video content like Instagram Reels and YouTube Shorts. The success of recovery narratives suggests that brief, personal videos may be particularly effective for health communication. Rather than focusing solely on restricting content, platforms might consider ways to actively promote recovery and harm reduction content, given its demonstrated appeal to users.
The emergence of TikTok as a recovery support space (Russell et al. 2022) offers valuable lessons for content moderation and platform design. While current approaches appear to effectively limit harmful content while allowing space for support and advocacy, refinements could address the challenges of moderating humor-based content. As short-form video becomes an increasingly dominant mode of social media communication, particularly among younger users, understanding how to foster supportive recovery communities while mitigating potential risks becomes important. Several areas warrant further research. While we found evidence of TikTok serving as a positive space for recovery support, there is a pressing need to understand the real-world impacts of this online discourse. Future studies should examine how exposure to recovery content affects help-seeking behaviors and treatment outcomes. Additionally, researchers should investigate potential risks and benefits for people who use drugs who engage with or create TikTok content. This is particularly important for content creators who publicly identify as being in active recovery, as sharing personal narratives may have complex psychological and social implications that are not yet well understood (Rennick-Egglestone et al. ). Additionally, investigating the intersection of online and offline support networks could provide insights into how digital recovery communities complement traditional recovery support services, optimizing online and in-person interventions. This work would be particularly valuable for reaching younger populations, who may be more likely to seek initial support through social media platforms.
Conclusion
This study explored the use of over 2.3K substance-related hashtags used on the popular short-form social media platform, TikTok. Through a semi-automatic community detection process, we identified 16 distinct hashtag communities, including Advocacy and Awareness, Health Conditions, and Humor. Our social network analysis revealed a highly interconnected small-world network (clustering coefficient 0.52, average path length 2.20) with influential hashtags like “#addiction”, “#recovery”, and “#sober” serving as bridges between communities. Through manual coding of 351 representative videos, we found that Recovery Advocacy content dominated (33.9% of videos), followed by Satirical content (28.2%), suggesting TikTok serves primarily as a platform for support rather than substance promotion. This is further supported by our finding that substances were present in only 91 videos (26%), with active use shown in just 23 videos. The alignment between hashtag communities and video content, particularly in recovery-focused videos, indicates organic community formation rather than attempts to evade content moderation. These findings have important implications for platform policies and interventions - while the prevalence of recovery-focused content suggests TikTok may serve as a valuable space for support and advocacy, the significant presence of humorous content alongside serious advocacy highlights the need for nuanced content moderation approaches. Future research should examine how these digital recovery communities complement traditional recovery support services and their real-world impacts on help-seeking behaviors and treatment outcomes, particularly among youth.
Limitations
@@ Conclusion
Several limitations of our study warrant discussion. First, our findings may not fully represent the entire TikTok ecosystem, as they are influenced by limitations in data availability and collection processes. TikTok’s (unofficial) API serves content through its recommendation algorithm, which means we received videos that did not necessarily contain our queried hashtags. While we attempted to address this by sampling videos across engagement quartiles to ensure representativeness, the algorithm’s opacity makes it difficult to guarantee complete coverage of substance-related content. This algorithmic bias may have also led to over-representation of certain creators whose content is favored by TikTok’s recommendation system. TikTok, in our experience, also cuts off data requests arbitrarily. Our approach of using 1,000 videos per hashtag was due to an alignment with the average number of videos we could reliably retrieve before being cut off on that request. Second, the dynamic nature of social media data presented challenges - some videos included in our initial network analysis were later unavailable for content coding due to accounts being banned or made private, a particular issue in communities discussing sensitive topics like substance use. Third, methodological challenges arose around identifying substance-related content, as users often employ common words as coded references to substances, making it difficult to definitively categorize some content as substance-related versus general discussion. This challenge was compounded by the platform’s rapid evolution of slang and terminology. Additionally, some videos contained duplicate hashtags, which could affect network metrics, and less popular hashtags may be underrepresented in our sample despite efforts to ensure broad coverage. These limitations highlight the challenges of studying substance-related content on social media platforms and suggest the need for complementary research approaches to fully understand these communities.
Appendix A Appendix
Appendix B Full Hashtag Network with Communities
@@
Figure 6 shows a visualization of the complete hashtag co-occurrence network, containing 2,333 nodes (hashtags) and 46,990 edges. The nodes are colored according to their community membership as determined by our semi-automated community detection process. The network layout was generated using a Fruchterman-Reingold algorithm, which positions nodes with stronger connections closer together. The central position of recovery-focused hashtags like “#addiction”, “#recovery”, and “#harmreduction” is clearly visible, demonstrating their role as bridges between communities. Node sizes are scaled according to their degree centrality, highlighting the most connected hashtags in the network. This visualization helps illustrate both the small-world properties of the network (clustering coefficient 0.52, average path length 2.20) and the organic formation of thematic communities around substance-related discourse on TikTok.
@@
Figure 6: The overall network of 2,333 hashtags related to substance use with colors representing each community as shown in the legend
Appendix C Network Centrality Measures per Community

### v2

@@ 4.3. RQ3: Effects of Substance Visibility
Given the observed patterns in how content creators use hashtags to frame and contextualize their videos, we next examined how the visibility of substances or related paraphernalia within videos relates to content type and engagement patterns. Videos were categorized into three groups based on substance presence: no substances shown ($n=260$), substances shown but not actively used ($n=68$), and active substance use ($n=23$). Videos showing active substance use were predominantly focused on documentation for social media (39.1% of such videos; Figure 5), with a stronger tendency towards lifestyle or vlog-style content. In contrast, recovery advocacy content was most prevalent in videos without substance visibility (38.8%), while promotional content appeared most frequently in videos that showed but did not depict active use of substances (29.4%; e.g., videos promoting nicotine products). Within videos showing substances, content predominantly featured alcohol, nicotine, or cannabis products, with only one video collage containing “hard drugs” such as Percocet, Adderall, and Xanax.
Kruskal-Wallis H tests (kruskal_use_1952) compared engagement metrics (views, likes, comments, and shares) across substance visibility categories (Table 4). Results indicated a statistically significant difference in likes ($H=6.97$, $p=.031$), with videos depicting active substance use showing higher median likes (Mdn =13,000=13{,}000) compared to videos showing substances without use (Mdn =982=982) or no substances (Mdn =1,028=1{,}028). Differences in views ($p=.087$), comments ($p=.171$), and shares ($p=.093$) did not reach statistical significance at $\alpha=.05$. However, descriptive patterns consistently showed higher median engagement for videos depicting active substance use across all metrics - notably, median views were over five times higher (79,100 vs. 14,900) and median comments nearly four times higher (150 vs. 39) compared to videos without substance visibility. Given the small sample size for active use videos ($n=23$), these results should be interpreted cautiously. Nevertheless, the findings suggest that the nature of substance depiction may influence audience engagement, with active use content attracting greater interaction despite comprising a small fraction of the overall substance-related content landscape.
Median (IQR)
Metric No Showing Using H p
$(n=260)$ $(n=68)$ $(n=23)$
Views 14,900 22,150 79,100 4.88 .087
(1,972–86,775) (968–234,175) (12,450–263,200)
Likes 1,028 982 13,000 6.97 .031*
(160–6,146) (43–13,275) (1,222–42,250)
Comments 39 34 150 3.53 .171
(7–173) (2–172) (21–253)
Shares 20 24 112 4.76 .093
(2–198) (1–458) (28–350)
Table 4. Engagement metrics by substance visibility category. Kruskal-Wallis H tests compared engagement across visibility levels. IQR = Interquartile Range. *$p<.05$
5. Discussion
Our findings present a critical case study of a socio-technical ecosystem where user agency and platform governance collide. While platforms like TikTok are often critiqued for amplifying harmful behaviors, including eating disorders (Bickham2024HiddenInPlainSight), self-harm content (lookingbill2024always), and dangerous viral challenges (Curtis et al., 2018; rutherford_turnttrending_2023), our analysis of its algorithmically surfaced content reveals a more complex reality. These insights emerge from a mixed-methods analysis that combines hashtag network structure with qualitative coding of algorithmically surfaced video content, allowing us to examine not only what content exists but what is made visible at scale. We find that, far from being a simple vector for promotion, the visible discourse on substance use is dominated by digital recovery narratives and harm reduction advocacy. This suggests that a combination of community-driven needs, user-led identity work, and active platform moderation has cultivated a powerful, large-scale recovery support space. This creates a fundamental tension: while the platform’s recommendation algorithm has documented harms, this same algorithmic infrastructure enables the formation of robust recovery support communities. This “algorithmic ambivalence” (phillips2018ambivalence) poses critical challenges for platform governance, as interventions designed to mitigate harm risk dismantling these beneficial support structures.
5.1. The Politics of Visibility in Digital Recovery Narratives
The dominance of recovery-focused content (33.9% of videos) suggests that TikTok functions as a venue for what Slade et al. (2021) term “digital recovery narratives.” These narratives serve a dual function. For creators, the act of public declaration serves as a form of ‘stigma management’ (goffman2009stigma), where the platform’s affordances for visibility are used to reframe a marginalized identity into one of strength and advocacy. For viewers, these narratives offer realistic models of recovery success, offer avenues for connection (Phelan2025SocialMediaRecovery), and provide ‘social proof’ (cialdini2004social) of sobriety, countering fatalistic views of addiction (corvini_impact_2024). This mirrors findings by li_explaining_2025, who identified that ideological homogeneity and central agendas are key to sustained participation in online movements; here, the ‘movement’ is recovery, and the shared identity is the driver of community.
However, this visibility raises questions regarding representation and power. While our findings demonstrate that recovery advocacy content achieves algorithmic prominence, understanding which recovery narratives are amplified, and which are marginalized, requires further investigation. The Phelan2025SocialMediaRecovery finding that individuals in SUD recovery actively use platforms like TikTok for support underscores the stakes of these visibility dynamics (Phelan2025SocialMediaRecovery). The absence of active use content in our sample does not necessarily imply its absence from the platform, but rather its algorithmic suppression or relegation to private spaces (gillespie2018custodians). This reflects a form of “uneven data shadows” (shelton2014mapping), where certain users and narratives, potentially including those in active addiction, those from marginalized communities, or those whose experiences don’t align with platform-favored content formats, remain algorithmically invisible. Consequently, claims about substance use on TikTok must be qualified as claims about what the platform’s governance systems allow to be widely visible.
5.2. Humor as Boundary Work and Stigma Management
The significant presence of humorous content (28.2%) alongside serious recovery narratives reflects a distinct form of health communication that we conceptualize as “boundary work” (gieryn1983boundary). In this context, humor helps users navigate the boundaries of platform acceptability and social stigma. By framing substance use experiences through satire and memes (shifman2014memes), creators may discuss addiction without the immediate vulnerability of serious or sensitive disclosure (andalibi2017sensitive), potentially lowering the barrier to entry for community participation. This finding aligns with rifat_combating_2024’s (rifat_combating_2024) observation that community-specific cultural practices—in their case, religious traditions, and in ours, vernacular humor—are essential for mediating online conflict and harm.
While research has shown that comedy can help individuals process difficult experiences (Martin, 2002), the prevalence of humor in this domain can also introduce risks of minimization. The algorithmic favoring of high-engagement humorous content (bucher2017algorithmic) could pressure creators to frame serious health crises comedically to achieve visibility, potentially distorting the perceived severity of substance use disorders for outside observers (chou2018addressing). This tension between therapeutic humor and potential minimization represents a key challenge for platforms hosting health-related content, suggesting that engagement metrics alone are insufficient proxies for the quality of health support.
5.3. Network Structure and the Cooperative Negotiation of Space
Our network analysis challenges the prevailing narrative of “adversarial algospeak,” where users are assumed to be constantly evading moderation through coded language (chancellor2016thyghgapp; Klug et al., 2023). While we observed some coded hashtags (e.g., #dr00gtiktok), the strong alignment between content types and straightforward hashtag communities (e.g., #harmreduction, #wedorecover) suggests that users are largely creating discourse within the bounds of platform affordances (boyd2007social), rather than purely against them. This indicates a cooperative, albeit fragile, negotiation of space where the community has successfully signaled the legitimacy of recovery discourse to the platform’s moderation systems, whether through community signaling, policy evolution, or emergent algorithmic behavior.
Structurally, we found that recovery-focused hashtags function as essential “bridge nodes” (burt1992structural) in the network, connecting otherwise siloed communities. This small-world (watts_collective_1998) property enables efficient information flow between disparate groups, such as those focused on specific substances and those focused on sobriety. This structural insight is crucial for governance: blunt-force moderation strategies, such as banning specific keywords, could sever these bridges, fragmenting the network and isolating users in riskier sub-communities without pathways to support (chandrasekharan2017you).
5.4. Implications for Harm-Aware Platform Design
Our findings suggest that platform governance must move beyond a binary of “remove” or “allow” (gillespie2018custodians) toward a strategy of context-sensitive amplification. We propose several concrete design directions:
1- Leverage network topology for harm reduction pathways. Our identification of recovery-focused hashtags as bridge nodes (#addiction, #recovery, #harmreduction) reveals structural leverage points for intervention. Rather than simply suppressing substance-related content, recommendation algorithms (resnick1997recommender) could proactively surface content from these bridge communities to users engaging with substance-specific or active-use clusters. For example, users repeatedly engaging with #opiates or #pingtok content could receive algorithmically-inserted suggestions for #harmreduction or #wedorecover content, creating pathways from substance-specific communities toward support resources.This approach aligns with public health principles of “meeting people where they are”(marlatt2012harm) while leveraging the small-world network structure (watts_collective_1998) we identified.
2- Differentiate content types in moderation systems. Our codebook (Table 3) demonstrates that recovery advocacy, harm reduction information, active use depiction, and community humor are distinguishable categories.Current moderation systems often treat all substance-related content uniformly (gillespie2018custodians), but context-sensitive classification—potentially incorporating community input (jhaver_does_2019) or subject-matter expertise—could enable more nuanced interventions. Content depicting active use without educational framing might be age-gated or de-amplified, while recovery narratives and harm reduction information could be preserved or even amplified. However, such systems risk creating new forms of surveillance (eubanks2025automating) and must be designed with strong privacy protections and appeal mechanisms (klonick2018new).
3- Design for semi-private peer support. The current “public-by-default” architecture of TikTok (INPLP_TikTok_2024) creates tensions between visibility (for community-building) and privacy (to avoid stigma (goffman2009stigma)). Platforms could offer middle-ground affordances: opt-in recovery communities visible to interested users but not broadcast to followers, pseudonymous accounts specifically for health disclosure, or fine-grained content controls allowing users to share recovery content with support networks while keeping it separate from professional/family audiences (marwick2011tweet). Such features would need careful design to avoid inadvertently marking recovery content as shameful.
4- Involve affected communities in governance. Our finding that hashtag communities show organic alignment rather than adversarial evasion suggests potential for collaborative governance (suzor2019we). Rather than purely top-down moderation, platforms could establish advisory structures including people with lived experience, harm reduction advocates, and public health professionals. Such approaches could help navigate difficult trade-offs between protecting users and preserving support communities, while ensuring governance decisions don’t reproduce existing power inequities in who gets to define “appropriate” recovery discourse (noble2018algorithms).
Critical to all these recommendations is recognition that the same affordances enabling harm also enable help. Well-intentioned changes risk unintended consequences (friedman2013value): special treatment could stigmatize recovery content, surveillance infrastructures built for safety could be weaponized against marginalized users (eubanks2025automating), and clinical norms imposed on peer communities could undermine their organic character (solomon2004peer). Effective governance requires explicitly designing for this duality rather than seeking one-sided solutions.
5.5. The Politics of Algorithmic Invisibility
Our methodology, analyzing algorithmically-served, publicly visible content, captures what TikTok’s governance systems allow to surface (gillespie2018custodians), not the totality of substance-related discourse. This is not merely a limitation but a finding: the dominance of recovery content in our sample reflects active platform curation (gillespie2018custodians) through moderation policies and algorithmic amplification.
What remains invisible reveals as much about platform governance as what is visible (shelton2014mapping). Content in private accounts, shadow-banned videos (gillespie2010politics), age-gated material, and algorithmically suppressed posts likely contain different narratives—potentially including active use content, harm reduction information deemed “controversial,” or experiences from marginalized users whose content may not align with algorithmically-favored formats or whose access to platform visibility is constrained. The absence of active use depiction in our sample (only 6.5%) does not prove TikTok is free of such content (bishop2018anxiety); rather, it demonstrates that TikTok’s recommendation system successfully prevents this content from wide circulation.
This raises critical questions: Should platforms amplify certain recovery modalities over harm reduction approaches? Who decides what constitutes “appropriate” substance use discourse? Our data suggests TikTok has already made these decisions, whether through explicit policy or emergent algorithmic behavior (gillespie2010politics), in ways that privilege certain recovery modalities and demographic groups. The most vulnerable users—those in active addiction, those with intersecting marginalized identities (crenshaw1991mapping), those in acute crisis—may be least likely to appear in algorithmically-surfaced public content (noble2018algorithms), raising concerns about whose needs platform governance serves.
6. Limitations
@@
Our findings must be interpreted within several important constraints. First, our findings may not fully represent the entire TikTok ecosystem, as they are influenced by limitations in data availability and collection processes. TikTok’s (unofficial) API serves content through its recommendation algorithm, which means we received videos that did not necessarily contain our queried hashtags. While we attempted to address this by sampling videos across engagement quartiles to ensure representativeness, the algorithm’s opacity makes it difficult to guarantee complete coverage of substance-related content. This algorithmic bias may have also led to over-representation of certain creators whose content is favored by TikTok’s recommendation system. TikTok, in our experience, also cuts off data requests arbitrarily. Our approach of using 1,000 videos per hashtag was due to an alignment with the average number of videos we could reliably retrieve before being cut off on that request. Second, the dynamic nature of social media data presented challenges - some videos included in our initial network analysis were later unavailable for content coding due to accounts being banned or made private, a particular issue in communities discussing sensitive topics like substance use. Third, methodological challenges arose around identifying substance-related content, as users often employ common words as coded references to substances, making it difficult to definitively categorize some content as substance-related versus general discussion. This challenge was compounded by the platform’s rapid evolution of slang and terminology. Additionally, some videos contained duplicate hashtags, which could affect network metrics, and less popular hashtags may be underrepresented in our sample despite efforts to ensure broad coverage. Lastly, our study analyzed content characteristics rather than user impacts. We cannot make claims about whether exposure to recovery content affects help-seeking behaviors, treatment outcomes, or relapse risks. The relationship between what people see on TikTok and their subsequent actions remains an open question. These limitations underscore a fundamental point: platform research always captures a mediated reality shaped by technical systems, corporate governance, and social dynamics. Claims about “what happens on TikTok” must be qualified as “what TikTok’s governance systems allow to be widely visible.”
7. Conclusion and Future Direction
This study challenges narratives about social media platforms and substance use. Through analysis of 2,333 hashtags and 351 videos, we found that TikTok’s algorithmically surfaced, publicly visible substance use content is dominated by recovery advocacy (33.9%) and humor-based coping (28.2%), not glamorization or promotion. Our network analysis revealed a highly interconnected small-world structure (clustering coefficient 0.52, average path length 2.20) where recovery-focused hashtags like #addiction, #recovery, and #sober function as critical bridges between thematic communities. Through manual coding of 351 representative videos, we found that Recovery Advocacy content dominated (33.9% of videos), followed by Satirical content (28.2%), while direct substance depiction appeared in only 26% of videos, with active use shown in just 6.5%—suggesting that despite documented harms from algorithmic recommendation systems, these same systems can enable valuable peer support communities around stigmatized health conditions.
However, our findings also reveal critical tensions inherent in platform governance. The dominance of recovery content in our sample reflects TikTok’s active curation through moderation policies and algorithmic amplification—raising questions about whose narratives are privileged and what remains invisible. The alignment between hashtag communities and content types indicates organic community formation rather than adversarial evasion, suggesting users have successfully negotiated space within platform affordances. Yet this cooperative relationship remains fragile: changes in moderation policy, algorithmic updates, or increased regulatory pressure could easily disrupt these communities, pushing discourse underground or fragmenting support networks.
Our findings open several critical avenues for future research. While we found evidence of TikTok serving as a positive space for recovery support, there is a pressing need to understand the real-world impacts of this online discourse. Future studies should examine how exposure to recovery content affects help-seeking behaviors and treatment outcomes. Additionally, researchers should investigate potential risks and benefits for people who use drugs who engage with or create TikTok content. This is particularly important for content creators who publicly identify as being in active recovery, as sharing personal narratives may have complex psychological and social implications that are not yet well understood (rennick-egglestoneMentalHealthRecovery2019). Additionally, investigating the intersection of online and offline support networks could provide insights into how digital recovery communities complement traditional recovery support services, optimizing online and in-person interventions. This work would be particularly valuable for reaching younger populations, who may be more likely to seek initial support through social media platforms.
8. Ethics Statement
This research was deemed exempt by our institution’s IRB as it analyzes publicly available data with no interaction with human subjects. In our analysis of TikTok videos, we followed established practices for ethical social media research (fiesler2020no) and we took several steps to protect user privacy: (1) all analysis was conducted on aggregated data, (2) no usernames or personally identifiable information are included in our results, and (3) our public dataset will only contain video URLs and associated metadata, allowing other researchers to hydrate the data while respecting user privacy. All collected and annotated data was stored on secure servers approved by our institution. These steps ensure that potential negative outcomes due to the use of these data are minimized while maintaining the reproducibility of our research. The dataset and associated code will be made available upon acceptance.
Appendix A.1 Full Hashtag Network with Communities
@@
Figure A.1 shows a visualization of the complete hashtag co-occurrence network, containing 2,333 nodes (hashtags) and 46,990 edges. The nodes are colored according to their community membership as determined by our semi-automated community detection process. The network layout was generated using a Fruchterman-Reingold algorithm, which positions nodes with stronger connections closer together. The central position of recovery-focused hashtags like #addiction, #recovery, and #harmreduction is clearly visible, demonstrating their role as bridges between communities. Node sizes are scaled according to their degree centrality, highlighting the most connected hashtags in the network. This visualization helps illustrate both the small-world properties of the network (clustering coefficient 0.52, average path length 2.20) and the organic formation of thematic communities around substance-related discourse on TikTok.
@@
Figure A.1. The overall network of 2,333 hashtags related to substance use with colors representing each community as shown in the legend
Appendix A.2 Network Centrality Measures per Community